OpenStudy (anonymous):

evaluate the integral ∫cosxcos2x

OpenStudy (lgbasallote):

let \[\cos (2x) = \cos^2 x - \sin^2 x\]

OpenStudy (anonymous):

cos x cos 2x= 1/2 (cos x+cos 3x) this will be easier to evaluate

OpenStudy (anonymous):

1/2 (sin x+ (sin3x/3))+c